# Victoria, BC, Canada

### v3

@@ Abstract
Mixed methods Research methods Methodology Guiding Principles
@@ 1 Introduction
Through this paper, we aim to demystify, clarify, and define MMR and its various properties, designs, and application scenarios for the SE research community. We encourage and call for SE researchers to use mixed methods and to do so in a more principled way. Our hope is that it will serve as the start of a robust conversation in the community about intentional, rigorous, and innovative use of MMR in SE, one that inevitably others will build upon in the future.
The rest of this paper is organized as follows. In Section 2, we present a brief background of the most prominent and relevant influences in mixed methods research design over the decades. In Section 3, we present four core principles that underlie mixed method research: methodological rationale, novel integrated insights, procedural rigor and ethical research. In Section 4, we discuss some notable properties of mixed method designs, as well as common mixed methods research designs. In Section 5, we present four scenarios of research studies that fictional researchers pursue. We describe the methodological choices these characters make and how they followed the four core principles mentioned in Section 3. In Section 6, we use the concept of antipatterns to present what not to do in mixed method research. In Section 7, we discuss mixed methods research designs with other research designs (such as multi-method, case studies and secondary studies) and share a call to action before concluding the paper in Section 8. This paper includes an Appendix which contains a glossary of terms we use throughout the paper for ease of reference.
@@ 2 A Brief Overview of Mixed Methods Research
It is only in the mid-20th century, with roots in disciplines such as sociology and anthropology, that mixed methods research (MMR) gained traction with the formalization of approaches by authors including Creswell and Plano Clark creswell_designing_2018 . According to these authors, “The formative period of mixed methods began in the 1950s and continued up until the 1980s. This period saw the initial interest in using more than one method in a study” but these authors did not explicitly identify these approaches as mixed methods. Furthermore, the different typologies of mixed method designs did not emerge until the 1990s maxwell_expanding_2016 ; Greene1989 . In the social sciences field, Campbell and Fiske’s paper from 1959 campbell_convergent_1959 is recognized as the first adoption of multiple (that included mixed) research methods for validation purposes. They defined the multiple method design as multi-operationalism since the authors propose a multitrait, multimethod matrix that intersects quaL-quaN methods with a specific set of traits as a means for enforcing validity of research conclusions. In the mid 1960s, Webb et al. webb_unobtrusive_1966 introduced the concept of triangulation, followed by Denzin in the 1970s denzin_research_1978 who detailed how to triangulate using both within-methods triangulation and between-methods triangulation, emphasizing the latter as a more powerful means to overcome bias and as a way to pragmatically mix methods.
Although the combination of quaL-quaN methods has been used in practice for many years, the first evidence of a more formal use of mixed methods is identifiable only in the late 1980s Kuhn1977 , as discussed by Guest and Fleming guest_mixed_2015 and Johnson et al. Johnson2004 ,johnson_toward_2007 who point out how mixed methods research combines qualitative and quantitative approaches, forming a third paradigm that complements traditional methods to yield more comprehensive, balanced, and insightful results. The common denominator in these early conceptualizations of mixed methods is that MMR combines quaL-quaN elements. It is only recently that researchers have turned their attention to proposing typologies of mixed methods, i.e., which describe how to structure and design a MMR by combining quaL and quaN elements according to some criteria as stated by Teddlie and Tashakkori teddlie_foundations_2009 .
@@
Today, MMR is widely recognized and utilized across diverse fields such as psychology, education, public health, information systems and management, and these approaches offer a flexible and robust framework for addressing multifaceted research inquiries and advancing knowledge in complex socio-technical domains poth_innovation_2018 ; ladner_mixed_2019 . In software engineering, we can learn much from related disciplines, such as information systems venkatesh2013bridging ; venkatesh2016guidelines , that articulate principles and guidelines of how to conduct mixed methods research. The presented guidelines by Venkatesh et al. venkatesh2013bridging , building on top of the mixed methods research by Tashakkori and Creswel TashakkoriCreswell2008 , Teddlie and Tashakkori teddlie2003major ; teddlie_foundations_2009 , Creswell creswell_research_2003 among others, clearly articulate that “the decision to conduct mixed methods research should hinge on the research question, purpose, and context.”
With the increasing complexity of MMR designs, the problem of discussing and visually presenting complex designs to readers became apparent. Consequently, as noted by Creswell and Plano Clark (poth_sage_2023, , Chapter 2) when they revisited MMR designs twenty years after their important chapter was first published in 2003 tashakkori_handbook_2003 , the most recent studies decreased their focus on the quaN and quaL components and increased attention towards the integration processes across design types. To reflect this change about the central elements of the MMR designs, Figure 1 represents the simplification of the interconnection of the core components of a MMR design. From this perspective, treating the data beyond quaL and quaN analysis and gaining insight through integration and meta-inferences [^2] See Appendix for a glossary of terms used through the paper. Integration is discussed in detail later in the paper. become central.
@@ 3 Principles to Guide Mixed Method Research
Two of the principles presented in our paper, methodological rationale and development of novel integrated insights (meta-inferences), relate to understanding why we should consider using a MMR design. The other two, procedural rigor and ethical research, relate to understanding how we should go about designing and applying MMR. While some aspects of these principles apply to single method applications and can be considered good research practice in general, others are particularly relevant to MMR and generally more complex in that context. We delve into these four principles below.
@@ 3.1 Methodological Rationale
Principle of Methodological Rationale: Mixed method research study designs should explain how the mixed research design and data methods are relevant for addressing a given research problem. By adhering to this principle, an MMR study description can address the question why did we select a particular MMR research design and set of procedures?
There are several conceptual reasons for using mixed methods [^3] The reasons we list are inspired by or earlier proposed by Ladner ladner_mixed_2019 , Bryman bryman_research_2007 , Creswell and Plano Clark creswell_designing_2018 and McGrath mcgrath_methodology_1995 .. The rationale for the methodology may be one or more of the following:
@@
  • Increased design flexibility: implies researchers may need to consider research design alternatives to offset limitations of a single method when faced with unexpected issues or unsatisfactory data. For example, many researchers that were planning to conduct ethnographic studies during the global pandemic pivoted to using alternative designs, such as a combination of online interviews, self-reported work diary entries, and data mining [^4] https://www.microsoft.com/en-us/research/project/the-new-future-of-work/.
@@ 3.2 Novel Integrated Insights
An MMR study should deliver on its promise of additional benefits to what could have been afforded by using a single method poth_innovation_2018 . Typically, once data analysis for each quaL/quaN method is complete, the researcher compiles the results and findings, (a phase often referred to as the inferential stage teddlie_foundations_2009 ). In an MMR study, inferences are also drawn from both quaL and quaN analyses. However, to achieve a truly integrated approach, these conclusions should be synthesized into meta-inferences (see Fig. 1), which may involve generating additional insights or linking findings from both methods to broader theories or implications [^5] Venkatesh et al. venkatesh2013bridging provide an in-depth discussion on development and assessment of the quality of meta-inferences.. An MMR study may bring novel insights in terms of one or more of the following:
@@
  • Complementary storytelling: quaN data provides a numerical strength and statistical measure, and quaL data provides richer insights into the experiences and perspectives of those close to the phenomenon. Together they help narrate a better and fuller story, with multiple ways to present findings in engaging formats [^6] Bryman refers to this as illustration, or the idea of “putting ‘meat on the bones’ of ‘dry’ quaN findings” bryman_social_2012 .
@@ 3.3 Procedural Rigor
  • Are the methods effectively integrated to answer the research question? It is important to appraise how well the data from the multiple methods are integrated to answer the research question posed and to appraise at what stage the methods are integrated: during collection or during analysis.
@@
  • Are the different methods used rigorously conducted? MMR studies should adhere to the rigor of the individual studies used in the mixed study (referring to the appropriate quality criteria for each study [^7] See the ACM SIGSOFT Empirical Standards for Software Engineering sigsoft . Other references could be consulted for further considerations on the aspects of quality and validation in the MMR integration process, such as Teddlie and Tashakkori teddlie_foundations_2009 and Venkatesh et al. venkatesh2013bridging ; venkatesh2016guidelines .).
@@ 3.4 Ethical Research
Principle of Ethical Research: Mixed methods research studies should manifest the application of responsible research, with respect for people (e.g., participants, concerned parties, beneficiaries) and their welfare (e.g., respect for their confidentiality, privacy, time, emotions, cultural sensitivity, and needs), and that of the environment (e.g., organisations, communities of practice, society, and the planet at large, aiming for sustainability). By adhering to this principle, an MMR study report can address the question how responsibly did we do it?
Ethical research is a timeless concept, with a push by institutions and publishing organizations towards the use of Ethics’ Boards to review studies for ethical compliance. However, the more recent unchecked advancements in AI and unprecedented access to publicly available datasets has meant that SE advancements are progressing faster than ethical standards, principles, and regulations can keep up. Both SE industry and research communities are calling for action to take stock of where we are headed collectively in terms of ethical practice in the software industry. Industrial and academic researchers alike need to be held accountable for, be aware of, and proactively manage the ethical aspects of research.
@@
  • Considering the why: Just because we may have access to large datasets and the latest in computational advancements, should we rush into designing research that can be performed using them or about them? What are the driving motivations for conducting the research in the first place? What benefits may it bring for the community being studied and for a wider society? Similarly, just because we can design a survey to ask some relevant questions, should we be spamming practitioner mailing lists with requests to fill them out? For example, contacting individuals, contributors, or leaders privately without explicit permission has been deemed as unacceptable behaviour by GitHub [^8] https://github.com/community/community/blob/main/CODE_OF_CONDUCT.md.
@@ 4.1 Design Properties
  • Research Questions: MMR questions differ from purely quantitative questions (that typically focus on description, comparison, or relationships between variables) or qualitative questions (that are typically open-ended, evolving, and exploratory witgh the aim to uncover processes or experiences). The MMR question addresses what the researcher hopes to learn through the combination (integration) of the quaL and quaN data creswell_designing_2018 . In turn, these questions will shape other design properties as listed next.
@@ 4.3 Research Designs
  • Explanatory Sequential: This sequential design starts with and prioritizes a deductive quaN method (normally the dominant method) that is followed by a quaL method to help explain the insights from the first method. The design may be planned up front or the need for the second method may emerge, especially in light of some unexpected insights or where some insights need further interpretation or explanation. For example, Jiang et al. Jiang2024 mined open source deep learning projects to analyze the distribution of computer vision defects, followed by interviews conducted with leaders of a reengineering team to shed light on the challenges and practices that may have led to the defects.
@@
5 MMR Design Scenarios
@@ 5.4 Zara
Ali (he, him), is a postdoc working closely with a partner company studying the introduction of a new security tool to ensure compliance with secure development practices across a large software company. Ali and his partners recognize that introducing a tool that enforces changes in development work has complex and likely unexpected impacts, not just on the security and or quality of the authored code, but also on the developers’ perception of their productivity and experience. His research goal is to identify what impacts the tool may have (within the context of the developers’ work and environment), and using these initial insights, investigate if these perceptions are shared broadly by other developers across the company. Ali’s research question asks: How does a development tool that enforces compliance with secure development practices impact developer perceived productivity and their experience?
@@
Through a literature review, Ali finds a lack of research on developers’ perceived impacts of the specific type of security tool he is studying. He aims to investigate whether developers believe that the tool improves the quality of the code but also how it may impact other aspects of their development experience. Ali could begin by designing a mono-method survey, such as one based on a validated technology adoption model tam89 , and distributing it widely across the company. However, his primary goal is to explore inductively what impacts the new security compliance tool might have, focusing on both the anticipated and unanticipated aspects of developers’ experiences. Furthermore, Ali aims to assess whether these initial findings can be generalized to a broader population of developers within the company. Ali’s goals could not be achieved without an integrative thinking approach supported by MMR, which leads him to use an exploratory sequential mixed methods design.
First, Ali interviews (quaL method) about 20 developers to uncover these impacts. Next, based on the initial insights, he conducts a follow-up survey (quaN method) across a broader population of developers at the studied company. A survey that is faster to answer was part of his rationale for this design. See Figure 3 for an overview of Ali’s MMR approach.
@@
The interviews and survey were conducted following respective best practices [^10] We do not elaborate on the rigor of the individual methods used to focus our paper on the discussion of rigor that applies to the mixing of methods.. As mentioned above, the rationale for using these two methods was primarily to develop the survey and to triangulate and generalize the findings across a broader sample. Ali aimed to conduct both methods (sequentially) at the start of the research, but considers the qualitative insights as the dominant contribution in this exploratory sequential research design. Ali, was careful not to invite those interviewed to answer the survey, as doing so could influence their survey responses. Ali also took care to word survey questions to avoid bias with insights gained through the interviews. Ali assesses the overall findings from MMR, not from individual studies but as an integrated contribution. He discusses the potential threats to validity that may arise during data collection and analysis when compiling the final results report for his study. For instance, it could have been the case that the follow up survey led to divergent results about the abandonment of other tools or on productivity experiences. Such divergences would need further investigation.
@@
As Ali conducts both methods, he takes care to protect the identity of the team members he interviewed and surveyed. Ensuring this confidentiality was especially important given security is a key concern of the organization, and developers may not feel free to share concerns about the friction a compliance tool may introduce. Ali also was careful how nuanced insights from the interviews (e.g., about abandoned tools which may have been used by specific developers only) were later framed in a neutral ways so that the small group of interviewed developers could not be identified. Ali also took care not to send the survey to the interviewed participants, both to ensure the integrity of the data (supporting rigour), but also to be respectful of participant time. The design itself was also done with the careful use of developer time in mind.

### v4

@@ Abstract
Mixed methods Research methods Methodology Guiding Principles Guidelines
@@ 1 Introduction
Through this paper, we aim to demystify, clarify, and define MMR and its various properties, designs, and application scenarios for the SE research community. We encourage and call for SE researchers to use mixed methods and to do so in a more principled way. Our hope is that it will serve as the start of a robust conversation in the community about intentional, rigorous, and innovative use of MMR in SE, one that others inevitably will build upon in the future.
The rest of this paper is organized as follows. In Section 2, we present a brief background of the most prominent and relevant influences in mixed methods research design over the decades. In Section 3, we present four core principles that underlie mixed method research: methodological rationale, novel integrated insights, procedural rigor and ethical research. In Section 4, we discuss some notable properties of mixed method designs, as well as common mixed methods research designs. In Section 5, we present four scenarios of research studies that fictional researchers pursue. We describe the methodological choices these characters make and how they followed the four core principles described in Section 3. In Section 6, we use the concept of antipatterns to present what not to do in mixed method research. In Section 7, we discuss mixed methods research designs with other research designs (such as multi-method, case studies and secondary studies) and share a call to action before concluding the paper in Section 8. This paper includes an Appendix which contains a glossary of terms we use throughout the paper for ease of reference.
@@ 2 A Brief Overview of Mixed Methods Research
It is only in the mid-20th century, with roots in disciplines such as sociology and anthropology, that mixed methods research (MMR) gained traction with the formalization of approaches by authors including Creswell and Plano Clark creswell_designing_2018 . According to these authors, the early development of mixed methods began in the 1950s and extended through the 1980s, marking the initial interest in incorporating multiple methods within a single study. These authors did not explicitly identify these approaches as mixed methods. Furthermore, the different typologies of mixed method designs did not emerge until the 1990s maxwell_expanding_2016 ; Greene1989 . In the social sciences field, Campbell and Fiske’s paper from 1959 campbell_convergent_1959 is recognized as the first adoption of multiple (that included mixed) research methods for validation purposes. They defined the multiple method design as multi-operationalism since the authors propose a multitrait, multimethod matrix that intersects quaL-quaN methods with a specific set of traits as a means for enforcing validity of research conclusions. In the mid 1960s, Webb et al. webb_unobtrusive_1966 introduced the concept of triangulation, followed by Denzin in the 1970s denzin_research_1978 who detailed how to triangulate using both within-methods triangulation and between-methods triangulation, emphasizing the latter as a more powerful means to overcome bias and as a way to pragmatically mix methods.
Although the combination of quaL-quaN methods has been used in practice for many years, the first evidence of a more formal use of mixed methods is identifiable only in the late 1980s Kuhn1977 , as discussed by Guest and Fleming guest_mixed_2015 and Johnson et al. Johnson2004 ,johnson_toward_2007 who point out how mixed methods research combines qualitative and quantitative approaches, forming a third paradigm that complements traditional methods to yield more comprehensive, balanced, and insightful results. The common denominator in these early conceptualizations of mixed methods is that MMR combines quaL-quaN elements. It is only recently that researchers have turned their attention to proposing typologies of mixed methods which describe how to structure and design an MMR by combining quaL and quaN elements according to some criteria as stated by Teddlie and Tashakkori teddlie_foundations_2009 .
@@
Today, MMR is widely recognized and utilized across diverse fields such as psychology, education, public health, information systems and management, and these approaches offer a flexible and robust framework for addressing multifaceted research inquiries and advancing knowledge in complex socio-technical domains poth_innovation_2018 ; ladner_mixed_2019 . In software engineering, we can learn much from related disciplines, such as information systems venkatesh2013bridging ; venkatesh2016guidelines , that articulate principles and guidelines of how to conduct mixed methods research. The guidelines by Venkatesh et al. venkatesh2013bridging , building on top of the mixed methods research by Tashakkori and Creswel TashakkoriCreswell2008 , Teddlie and Tashakkori teddlie2003major ; teddlie_foundations_2009 , Creswell creswell_research_2003 among others, clearly state that the choice to conduct mixed methods research should depend on the research question, purpose, and context.
With the increasing complexity of MMR designs, the problem of discussing and visually presenting complex designs to readers became apparent. Consequently, as noted by Creswell and Plano Clark (poth_sage_2023, , Chapter 2) when they revisited MMR designs twenty years after their important chapter was first published in 2003 tashakkori_handbook_2003 , the most recent studies decreased their focus on the quaN and quaL components and increased attention towards the integration processes across design types. To reflect this change about the central elements of the MMR designs, Figure 1 represents the simplification of the interconnection of the core components of an MMR design. From this perspective, treating the data beyond quaL and quaN analysis and gaining insight through integration and meta-inferences [^2] See Appendix for a glossary of terms used through the paper. Integration is discussed in detail later in the paper. become central.
@@ 3 Principles to Guide Mixed Method Research
Two of the principles presented in our paper, methodological rationale and development of novel integrated insights (meta-inferences), relate to understanding why we should consider using an MMR design. The other two, procedural rigor and ethical research, relate to understanding how we should go about designing and applying MMR. While some aspects of these principles apply to single method applications and can be considered good research practice in general, others are particularly relevant to MMR and generally more complex in that context. We delve into these four principles below.
@@ 3.1 Methodological Rationale
Principle of Methodological Rationale: Mixed method research study designs should explain how the mixed research design and data methods are relevant for addressing a given research problem. By adhering to this principle, an MMR study description can address the question why did we select a particular MMR design and set of procedures?
There are several conceptual reasons for using mixed methods. [^3] The reasons we list are inspired by or earlier proposed by Ladner ladner_mixed_2019 , Bryman bryman_research_2007 , Creswell and Plano Clark creswell_designing_2018 and McGrath mcgrath_methodology_1995 . The rationale for the methodology may be one or more of the following:
@@
  • Increased design flexibility: mixed methods may allow researchers to consider research design alternatives to offset limitations of a single method when faced with unexpected issues or unsatisfactory data. For example, many researchers that were planning to conduct ethnographic studies during the global pandemic pivoted to using alternative designs, such as a combination of online interviews, self-reported work diary entries, and data mining. [^4] https://www.microsoft.com/en-us/research/project/the-new-future-of-work/
@@ 3.2 Novel Integrated Insights
An MMR study should deliver on its promise of additional benefits to what could have been afforded by using a single method poth_innovation_2018 . Typically, once data analysis for each quaL/quaN method is complete, the researcher compiles the results and findings (a phase often referred to as the inferential stage teddlie_foundations_2009 ). In an MMR study, inferences are also drawn from both quaL and quaN analyses. However, to achieve a truly integrated approach, these conclusions should be synthesized into meta-inferences (see Fig. 1), which may involve generating additional insights or linking findings from both methods to broader theories or implications. [^5] Venkatesh et al. venkatesh2013bridging provide an in-depth discussion on development and assessment of the quality of meta-inferences. An MMR study may bring novel insights in terms of one or more of the following:
@@
  • Complementary storytelling: quaN data provides a numerical strength and statistical measure, and quaL data provides richer insights into the experiences and perspectives of those close to the phenomenon. Together they help narrate a better and fuller story, with multiple ways to present findings in engaging formats. [^6] Bryman refers to this as illustration, or the idea of “putting ‘meat on the bones’ of ‘dry’ quaN findings” bryman_social_2012
@@ 3.3 Procedural Rigor
  • Are the methods effectively integrated to answer the research question? It is important to appraise how well the data from the multiple methods are integrated to answer the research question posed and to appraise at what stage the methods are integrated: during collection, during analysis or during both collection and analysis.
@@
  • Are the different methods used rigorously conducted? MMR studies should adhere to the rigor of the individual studies used in the mixed study (referring to the appropriate quality criteria for each study). [^7] See the ACM SIGSOFT Empirical Standards for Software Engineering ralph2020empirical . Other references could be consulted for further considerations on the aspects of quality and validation in the MMR integration process, such as Teddlie and Tashakkori teddlie_foundations_2009 and Venkatesh et al. venkatesh2013bridging ; venkatesh2016guidelines .
@@ 3.4 Ethical Research
Principle of Ethical Research: Mixed methods research studies should manifest the application of responsible research, with respect for people (e.g., participants, concerned parties, beneficiaries) and their welfare (e.g., respect for their confidentiality, privacy, time, emotions, cultural sensitivity, and needs), and that of the environment (e.g., organizations, communities of practice, society, and the planet at large, aiming for sustainability). By adhering to this principle, an MMR study report can address the question how responsibly did we do it?
Ethical research is a timeless concept, with a push by institutions and publishing organizations towards the use of Ethics Boards to review studies for ethical compliance. However, the more recent unchecked advancements in AI and unprecedented access to publicly available datasets has meant that SE advancements are progressing faster than ethical standards, principles, and regulations can keep up. Both SE industry and research communities are calling for action to take stock of where we are headed collectively in terms of ethical practice in the software industry. Industrial and academic researchers alike need to be held accountable for, be aware of, and proactively manage the ethical aspects of research.
@@
  • Considering the why: Just because we may have access to large datasets and the latest in computational advancements, should we rush into designing research that can be performed using them or about them? What are the driving motivations for conducting the research in the first place? What benefits may it bring for the community being studied and for a wider society? Similarly, just because we can design a survey to ask some relevant questions, should we be spamming practitioner mailing lists with requests to fill them out? For example, contacting individuals, contributors, or leaders privately without explicit permission has been deemed as unacceptable behaviour by GitHub. [^8] https://github.com/community/community/blob/main/CODE_OF_CONDUCT.md
@@ 4.1 Design Properties
  • Research Questions: MMR questions differ from purely quantitative questions (that typically focus on description, comparison, or relationships between variables) or qualitative questions (that are typically open-ended, evolving, and exploratory with the aim to uncover processes or experiences). The MMR question addresses what the researcher hopes to learn through the combination (integration) of the quaL and quaN data creswell_designing_2018 . In turn, these questions will shape other design properties as listed next.
@@ 4.3 Research Designs
  • Explanatory Sequential: This sequential design starts with and prioritizes a deductive quaN method (normally the dominant method) that is followed by a quaL method to help explain the insights from the first method. The design may be planned up front or the need for the second method may emerge, especially in light of some unexpected insights or where some insights need further interpretation or explanation. For example, Jiang et al. Jiang2024 mined open source deep learning projects to analyze the distribution of computer vision defects, followed by interviews conducted with leaders of a reengineering team, to shed light on the challenges and practices that may have led to the defects.
@@
5 Mixed Method Research Design Scenarios
@@ 5.4 Zara
Ali (he, him), is a postdoc working closely with a partner company studying the introduction of a new security tool to ensure compliance with secure development practices across a large software company. Ali and his industry partner recognize that introducing a tool that enforces changes in development work has complex and likely unexpected impacts, not just on the security and or quality of the authored code, but also on the developers’ perception of their productivity and experience. His research goal is to identify what impacts the tool may have (within the context of the developers’ work and environment), and using these initial insights, investigate if these perceptions are shared broadly by other developers across the company. Ali’s research question asks: How does a development tool that enforces compliance with secure development practices impact developer perceived productivity and their experience?
@@
Through a literature review, Ali finds a lack of research on developers’ perceived impacts of the specific type of security tool he is studying. He aims to investigate whether developers believe that the tool improves the quality of the code but also how it may impact other aspects of their development experience. Ali could begin by designing a mono-method survey, such as one based on a validated technology adoption model tam89 , and distributing it widely across the company. However, his primary goal is to explore inductively what impacts the new security compliance tool might have, focusing on both the anticipated and unanticipated aspects of developers’ experiences. Furthermore, Ali aims to assess whether these initial findings can be generalized to a broader population of developers within the company. Ali’s goals could not be achieved without an integrative thinking approach supported by MMR, which leads him to use an exploratory sequential mixed methods design. First, Ali interviews (quaL method) about 20 developers to uncover these impacts. Next, based on the initial insights, he conducts a follow-up survey (quaN method) across a broader population of developers at the studied company. A survey that is faster to answer was part of his rationale for this design. See Figure 3 for an overview of Ali’s MMR approach.
@@
The interviews and survey were conducted following respective best practices. [^10] We do not elaborate on the rigor of the individual methods used to focus our paper on the discussion of rigor that applies to the mixing of methods. As mentioned above, the rationale for using these two methods was primarily to develop the survey and to triangulate and generalize the findings across a broader sample. Ali aimed to conduct both methods (sequentially) at the start of the research, but considers the qualitative insights as the dominant contribution in this exploratory sequential research design. Ali was careful not to invite those interviewed to answer the survey, as doing so could influence their survey responses. Ali also took care to word survey questions to avoid bias with insights gained through the interviews. Ali assesses the overall findings from MMR, not from individual studies but as an integrated contribution. He discusses the potential threats to validity that may arise during data collection and analysis when compiling the final results report for his study. For instance, it could have been the case that the follow up survey led to divergent results about the abandonment of other tools or on productivity experiences. Such divergences would need further investigation.
@@
As Ali conducts both methods, he takes care to protect the identity of the team members he interviewed and surveyed. Ensuring this confidentiality was especially important given security is a key concern of the organization, and developers may not feel free to share concerns about the friction a compliance tool may introduce. Ali was also careful how nuanced insights from the interviews (e.g., about abandoned tools which may have been used by specific developers only) were later framed in a neutral ways so that the small group of interviewed developers could not be identified. Ali also took care not to send the survey to the interviewed participants, both to ensure the integrity of the data (supporting rigor), but also to be respectful of participant time. The design itself was also done with the careful use of developer time in mind.